# Intergalactic magnetic field lower limits up to the redshift $z\approx 3$

### v1

@@ 3.2 IGMF volume-filling fraction
The resulting limit on $f$ as a function of IGMF is given in Fig. 3. Marginalized over the IGMF values, the minimal IGFM volume-filling fraction required for the model agreement with Fermi/LAT data is $f>0.96$ at 95% confidence level. A similarly strong constraint is obtained in the $z\in[1;2]$ redshift window, dominated by RBS 1432. Exclusion of this source relaxes the $z\in[1;2]$ limit to $f>0.87$, close to values derived in the other two redshift bins ($f>0.83$ and $f>0.78$ for $z\in[0.5;1]$ and $z\in[2;3]$ correspondingly).
@@ 4 Discussion
The derived here limits present the first attempt to constrain IGMF using a single-instrument observations of both intrinsic and “echo” emission. Compared to earlier works, this approach eliminates the need in artificial assumptions regarding the source (non-)variability [e.g. 17, 18, 19, 35, 21] or spectral shape stability [e.g. 20], representing the key practical uncertainties behind earlier IGMF searches.
@@
such measurements alone may be not suitable for nearby sources, where the majority of IGMF lower limits has been derived thus far. Likewise, higher-redshift Fermi/LAT measurements can constrain only the high-energy tail of the “echo” emission – even if in the attenuated flux would be accurately measured with the next-generation VHE facilities like CTA [53]. To boost the “echo” sensitivity, those would have to be paired with dedicated observations with MeV-band instruments [such as AMEGO, e-ASTROGAM or MAST; 54, 55, 56].
It should be noted, that the limits obtained here may be affected by the development of the plasma instabilities, suggested to be capable of dissipating the absorbed emission power before it is re-emitted in the form of the “echo” [e.g. 57]. While presently there is no self-consistent description of this problem, recent studies indicate only a marginal role of the instabilities in reduction of the emission power available to the “echo” [58, 59]. As such, the effect of these instabilities was not considered here.
The derived here IGMF limits are lower than those earlier obtained from AGN observations, as shown in Fig. 5. In part this is due the choice of the VHE attenuation model, selected to minimize the predicted “echo” power for a given source emission spectrum. While this choice reduces the IGMF exclusion power in this work, it makes the derived limits robust against existing uncertainties of the VHE attenuation. Furthermore, these limits provide the first direct evidence for the presence of IGMF at $z\gtrsim 1$, enabling tests of its cosmological evolution.
Figure 5: Comparison of the obtained here at 95% confidence level sample-averaged (blue) and individual-source (orange) comoving IGMF strength constraints for $\lambda_{B}=1$ Mpc with those derived earlier at lower redshifts. For sample-averaged constraints from literature [35, 20, 21, 60] the uncertainties indicate the redshift range of the sources employed; in case of GRB 221009A, a range of limits from Refs. [22, 23, 24] is shown.
@@
The galactic origin of comparably strong magnetic field on these spatial scales is disputed by the modelling of magnetogenesis around young galaxies [52, 7, 61, 28], suggesting the volume-filling fraction of $B>10^{-18}$ cG IGMF generated this way does not exceed $\sim 40\%$ even at $z\sim 0$. With the latter decreasing quickly above $z\sim 1$ [62, 7, 61, 28], the obtained here lower limits at $z\sim 2-3$ appear inconsistent with these calculations. Small volume-filling fraction of galactic-origin fields at $z\gtrsim 1$ implies the presented limits probe IGMF largely unaffected by the galactic feedback, pointing to its primordial origin. In this context it is worthy to point out that stronger IGMF limits at redshifts $z\sim 0$ do not necessarily constrain the same component of the present-day IGMF as the latter may still be partially contaminated by the galactic feedback. In this respect, IGMF measurements at high and low redshifts may be eventually used to differentiate the galactic and cosmological contributions to the present-day field. This possibility may be further clarified with further observations of with dedicated VHE instruments, such as LHAASO [63] and CTA [53] at low and high redshifts respectively.
@@ Appendix A EBL models comparison
Figure 6: Comparison of the energy at the optical depth $\tau=1$ for EBL models [29, 64, 65, 66, 49] supported in CRBeam and the newer one from Ref. [51]. Except for Ref. [49] model (brown curve), the energies were calculated with the ebltable package [67]. For models from Ref. [29] and [66], depicted by blue and red lines, values shown correspond to the redshift range covered by the tables in the original publications.
@@ Appendix B High-redshift “echo” signal spatial extension and time delay
CRBeam simulations enable direct estimates of the spatial extension and time delay of the “echo” signal in the presence of IGMF for the studied sources from their outputs, recording the final position and direction of all the “echo” photons as they reach the observer. The time delay was estimated as the source-to-observer and the finial source-to-photon distances difference divided by the speed of light; the spatial extension was estimated as a 68% quantile of the “echo” photons arrival direction offset with respect to the observer-to-source axis. The computed extension and time delay for selected sources, covering the redshift range of the AGNs in Tab. 1 are shown in Figs. 7 and 8 respectively, where they are compared with the Fermi/LAT point spread function and observational window duration. To account for the latter, expected “echo” extension was also recalculated using only photons arriving to the observer with the time delay less than 17 yr, as shown in Fig. 7 with dotted lines.
@@
Figure 7: Spatial extension (68% containment radius, solid lines) of the “echo” emission in the $E_{\gamma}=0.25-0.35$ GeV energy range, calculated for several representative sources from Tab. 1, covering the broad redshift range $z\approx 0.6-1.2$. Dashed lines of the same colors denote the “echo” extension if only time delays of $t_{d}<17$ yr, corresponding to the duration of Fermi/LAT observations, are considered. For comparison, the grey dashed line marks the Fermi/LAT angular resolution (same 68% containment) at 1 GeV photon energy.
Figure 8: Median time delay of the “echo” photons in the $E_{\gamma}=0.25-0.35$ GeV energy range for the same sources as in Fig7. For a given IGMF strength the calculated delay varies from source to source due to its strong $t_{d}\sim(1+z)^{-5}$ redshift dependence [15]. The horizontal dashed line marks the duration of the Fermi/LAT observations employed here.

### v2

@@ 3.2 IGMF volume-filling fraction
The resulting limit on $f$ as a function of IGMF is given in Fig. 3. Marginalized over the IGMF values, the minimal IGFM volume-filling fraction required for the model agreement with Fermi/LAT data is $f>0.96$ at 95% confidence level. A similarly strong constraint is obtained in the $z\in[1;2]$ redshift window, dominated by RBS 1432. Exclusion of this source relaxes the $z\in[1;2]$ limit to $f>0.87$, close to values derived in the other two redshift bins ($f>0.65$ and $f>0.78$ for $z\in[0.5;1]$ and $z\in[2;3]$ correspondingly).
@@ 4 Discussion
The derived here limits present the first attempt to constrain IGMF using single-instrument observations of both intrinsic and “echo” emission. Compared to earlier works, this approach eliminates the need in artificial assumptions regarding the source (non-)variability [e.g. 24, 63, 66, 6, 9] or spectral shape stability [e.g. 4], representing the key practical uncertainties behind earlier IGMF searches.
@@
such measurements alone may be not suitable for nearby sources, where the majority of IGMF lower limits has been derived thus far. Likewise, higher-redshift Fermi/LAT measurements can constrain only the high-energy tail of the “echo” emission – even if the attenuated flux would be accurately measured with the next-generation VHE facilities like CTA [7]. To boost the “echo” sensitivity, those would have to be paired with dedicated observations with MeV-band instruments [such as AMEGO, e-ASTROGAM or MAST; 47, 22, 28].
It should be noted, that the limits obtained here may be affected by the development of the plasma instabilities, suggested to be capable of dissipating the absorbed emission power before it is re-emitted in the form of the “echo” [e.g. 20]. While presently there is no self-consistent description of this problem, recent studies indicate only a marginal role of the instabilities in reduction of the emission power available to the “echo” [56, 13]. As such, the effect of these instabilities was not considered here.
The IGMF limits derived here are lower than those obtained earlier from AGN observations, as shown in Fig. 5. In part this is due the choice of the VHE attenuation model, selected to minimize the predicted “echo” power for a given source emission spectrum. While this choice reduces the IGMF exclusion power in this work, it makes the derived limits robust against existing uncertainties of the VHE attenuation. Furthermore, these limits provide the first direct evidence for the presence of IGMF at $z\gtrsim 1$, enabling tests of its cosmological evolution.
Figure 5: Comparison of the comoving IGMF strength constraints obtained here at 95% confidence level for $\lambda_{B}=1$ Mpc – both sample-averaged (blue) and for individual sources (orange) – with those derived earlier at lower redshifts. For sample-averaged constraints from literature [6, 4, 9, 19] the uncertainties indicate the redshift range of the sources employed; in case of GRB 221009A, a range of limits from Refs. [27, 36, 67] is shown.
@@
The galactic origin of comparably strong magnetic field on these spatial scales is disputed by the modelling of magnetogenesis around young galaxies [25, 32, 14, 64], suggesting the volume-filling fraction of $B>10^{-18}$ cG IGMF generated this way does not exceed $\sim 40\%$ even at $z\sim 0$. With the latter decreasing quickly above $z\sim 1$ [17, 32, 14, 64], the lower limits obtained here at $z\sim 2-3$ appear inconsistent with these calculations. Small volume-filling fraction of galactic-origin fields at $z\gtrsim 1$ implies the presented limits probe IGMF largely unaffected by the galactic feedback, pointing to its primordial origin. In this context it is worthy to point out that stronger IGMF limits at redshifts $z\sim 0$ do not necessarily constrain the same component of the present-day IGMF as the latter may still be partially contaminated by the galactic feedback. In this respect, IGMF measurements at high and low redshifts may be eventually used to differentiate the galactic and cosmological contributions to the present-day field. This possibility may be further clarified with further observations with dedicated VHE instruments, such as LHAASO [21] and CTA [7] at low and high redshifts respectively.
@@ Appendix A EBL models comparison
Figure 6: Comparison of the energy at the optical depth $\tau=1$ for EBL models [30, 40, 37, 31, 59, 61, 29] supported in CRBeam and the newer one from Ref. [59]. Except for Ref. [61] model (brown curve), the energies were calculated with the ebltable package [48]. For models from Ref. [30] and [31], depicted by blue and orange lines, values shown correspond to the redshift range covered by the tables in the original publications.
@@ Appendix B High-redshift “echo” signal spatial extension and time delay
CRBeam simulations enable direct estimates of the spatial extension and time delay of the “echo” signal in the presence of IGMF for the studied sources with their outputs recording the final position and direction of all the “echo” photons as they reach the observer. Based on them, the time delay was estimated here as the source-to-observer and the finial source-to-photon distances difference divided by the speed of light; the spatial extension was estimated as a 68% quantile of the “echo” photons arrival direction offset with respect to the observer-to-source axis. The computed extension and time delay for selected sources, covering the redshift range of the AGNs in Tab. 1 and roughly representing the employed redshift bin edges, are shown in Figs. 7 and 8 respectively, where they are compared with the Fermi/LAT point spread function and observational window duration. To account for the latter, expected “echo” extension was also recalculated using only photons arriving to the observer with the time delay less than 17 yr, as shown in Fig. 7 with dotted lines.
@@
Figure 7: Spatial extension (68% containment radius, solid lines) of the “echo” emission in the $E_{\gamma}=0.25-0.35$ GeV energy range, calculated for several representative sources from Tab. 1, covering the broad redshift range $z\approx 0.6-1.2$. Dashed lines of the same colors denote the “echo” extension if only time delays of $t_{d}<17$ yr, corresponding to the duration of Fermi/LAT observations, are considered. For comparison, the grey dashed line marks the Fermi/LAT angular resolution (same 68% containment, P8R3 SOURCE event type without sub-division into PSF types, acceptance weighted) at 1 GeV photon energy.
Figure 8: Median time delay of the “echo” photons in the $E_{\gamma}=0.25-0.35$ GeV energy range for the same sources as in Fig.7. For a given IGMF strength the calculated delay varies from source to source due to its strong $t_{d}\sim(1+z)^{-5}$ redshift dependence [52]. The horizontal dashed line marks the duration of the Fermi/LAT observations employed here.